Circular No. 7128

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999as and 1999at IN ANONYMOUS GALAXIES R. Knop, G. Aldering, S. Deustua, G. Goldhaber, M. Kim, and P. Nu-

gent, Lawrence Berkeley National Laboratory, on behalf of the Supernova Cosmology Project (SCP), report the discovery of two supernovae on Feb. 18 UT as part of a joint collaboration between the SCP and E. Helin, S. Pravdo, D. Rabinowitz, and K. Lawrence (Near-Earth Asteroid Tracking project, Jet Propulsion Laboratory):

SN 1999as is located at α = 9^{h}16^{m}30^{s}.86, δ = +13^{o}39′02′′.2 (equinox J2000.0), on the edge of a very faint galaxy. A CCD spectrum (range 430–890 nm) taken by Aldering at the MDM 2.4-m telescope on Mar. 6, shows the supernova to be an unusually blue object with broad absorption features. Further spectra, obtained Mar. 16 by A. Goobar and T. Dahlen (Stockholm University) and I. Hook (Royal Observatory, Edinburgh) using the Nordic Optical Telescope, and by Aldering using the Cerro Tololo Interamerican Observatory (CTIO) 4-m telescope, bear a resemblance to type-Ia supernovae, 1–2 weeks after maximum light, at an approximate redshift of 0.12, with the following caveats: several narrow spectral features consistent with Fe II lines are present; the Si II features are similar to those of SN 1991T; and a type-Ia supernova at this redshift should show a significant emission feature at 475 nm (yet these spectra show a large deficit of flux in this region). Finally, it should be noted that, at this redshift, the supernova is ∼ 2 mag brighter than a typical type-Ia supernova at maximum light. Currently SN 1999as is at V = 18.3.

SN 1999at (unfiltered mag 18.8) is located at α = 10^{h}23^{m}11^{s}.30, δ = +17^{o}59′06′′.0 (equinox 2000.0). An R-band image from the YALO 1-m telescope obtained Mar. 7 by D. G. Huerta (CTIO), A. Mourão (Centro de Astrofísica, Instituto Superior Tecnico, Lisbon), and Rui Agostinho, Joao Lin Yun, and F. D. Santos (University of Lisbon), confirmed the presence of the supernova and that the host was an elliptical galaxy. A CCD spectrum obtained at the ARC 3.5-m telescope on Mar. 10 by G. Richards (University of Chicago) and H. Newberg and S. Kent (Fermi National Accelerator Laboratory) shows the host at a redshift of 8180 km/s and a supernova consistent with a normal type-Ia event, < 10 days after maximum light. Finding charts for the supernovae can be obtained

from http://panisse.lbl.gov/nearsearch/fcharts/index.html. Su-

pernovae and comparison images can be found in the SkyMorph archive

at http://skys.gsfc.nasa.gov/skymorph/skymorph.html.

1999 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green